# Visual Autoregressive Modeling for Instruction-Guided Image Editing

### v1

@@ Introduction
Instruction-guided image editing (Brooks et al. 2023; Zhang et al. 2023; Sheynin et al. 2024) represents a significant advance in generative AI, shifting the paradigm from pure synthesis towards fine-grained interactive control. This progress has been propelled by two primary factors: the curation of large-scale, high-quality editing datasets (Ge et al. 2024; Ye et al. 2025) and the rise of diffusion models (Ho et al. 2020; Peebles et al. 2023; Labs 2024; Qian et al. 2025, 2024; Wan et al. 2025) as the mainstream architecture. In particular, the impressive visual fidelity is largely derived from the iterative denoising process in diffusion models. However, this core mechanism also imposes a critical limitation. The global nature of the denoising process often leads to unintended spurious edits, causing modifications to “bleed” into regions that should remain unchanged. Simultaneously, the multi-step denoising process incurs substantial computational cost and limits real-time practical use.
In contrast to diffusion models, autoregressive (AR) models (Esser et al. 2021; Sun et al. 2024; Pang et al. 2025) offer a fundamentally different paradigm. Instead of holistic iterative refinement, they synthesize an image causally in a sequential token-by-token manner. This compositional generation process is a natural fit for image editing. Specifically, it provides a flexible mechanism for preserving unchanged regions while precisely modifying edited regions, addressing the entanglement problem of diffusion models. More encouragingly, the pioneering visual autoregressive (VAR) modeling (Tian et al. 2024) breaks the spatial structural degradation through the scale-by-scale prediction strategy, leading to high-quality image generation. While the VAR paradigm has shown competitive performance in image synthesis (Kumbong et al. 2025; Han et al. 2025), its potential for instruction-guided editing remains largely unexplored. Early training-free attempts (Wang et al. 2025) repurpose text-to-image models but lack the task-specific knowledge required for robust editing, causing them to lag significantly behind diffusion-based methods (Hou et al. 2024; Zhao et al. 2024a; Zhang et al. 2025).
@@
Figure 1: VAREdit achieves high-precision performance in instruction-guided image editing. It excels across diverse and challenging editing scenarios, including object-level modifications (addition, replacement, removal), attribute changes (material, text, posture, style, color) and complex compositional edits.
@@
We address this critical gap by introducing VAREdit, the first tuning-based VAR framework for instruction-guided editing. VAREdit reframes image editing as a next-scale prediction problem, which generates multi-scale target features to achieve precise image modification. A central challenge in this paradigm is how to effectively condition the model on the source image. A straightforward approach is to present full-scale source conditions, yet it is computationally inefficient due to the lengthy input sequences. While the finest-scale-only conditional setting provides sufficient information and eliminates the redundant input tokens, it creates a severe scale mismatch since high-frequency details may disrupt the prediction of coarse target features. To resolve this issue, we perform a systematic analysis of inter-scale dependencies and uncover a key insight: the model’s sensitivity to scale-aware conditioning is concentrated in its first self-attention layer. Accordingly, we propose the Scale-Aligned Reference (SAR) module, a mechanism that injects the scale-matched source information only in the first self-attention layer, while keeping other layers still use the finest-scale condition. As demonstrated in Figure 1, this design enables VAREdit to achieve exceptionally precise editing.
Our contributions can be summarized as follows:
@@
  • We introduce VAREdit, the first tuning-based visual autoregressive model for instruction-guided image editing.
@@
  • We identify a scale mismatch issue in the finest-scale conditioned VAREdit and propose the SAR module as an effective solution.
@@
  • VAREdit sets a new state-of-the-art on standard editing benchmarks, outperforming leading diffusion models in both editing adherence and generation efficiency.
Related Work
Instruction-Guided Image Editing
@@
Instruction-guided image editing (Brooks et al. 2023; Zhao et al. 2024a; Yu et al. 2025; Cai et al. 2025) seeks to modify an image according to user’s textual instruction, which differs from caption-based approaches (Hertz et al. 2022; Feng et al. 2025) that often resynthesize an entire image from a global description. This field is predominantly driven by diffusion models and seminal work InstructPix2Pix (Brooks et al. 2023) established a powerful paradigm by channel-wise concatenating source and target images. Subsequent research advanced this by curating larger-scale datasets with “expert” models (Zhang et al. 2023; Sheynin et al. 2024; Zhao et al. 2024a; Ge et al. 2024; Ye et al. 2025) or exploring alternative conditioning, like spatial concatenation for large-motion edits (Zhang et al. 2025; Cai et al. 2025). However, diffusion-based models often produce unintended modifications and suffer from slow iterative sampling. To address these limitations, autoregressive (AR) models have been explored. For instance, EditAR (Mu et al. 2025) presents a vanilla AR editing framework, yet its performance is not on par with leading diffusion methods due to the spatial dependency limitation in base model. Our work builds on the pioneering visual autoregressive (VAR) (Tian et al. 2024) architecture, a tuning-based approach that surpasses recent diffusion-based models in both edit quality and efficiency.
Autoregressive Image Generation
@@
Autoregressive (AR) models, a mainstay in language tasks, have been extended to image synthesis. They operate by tokenizing an image into a discrete sequence and then training a Transformer model to predict the next token. Early works use vector quantization (VQ) for tokenization (Van Den Oord et al. 2017; Esser et al. 2021; Chang et al. 2022; Yu et al. 2022), while recent advances such as lookup-free quantization (Yu et al. 2023; Zhao et al. 2024b) improve performance with large vocabularies. The AR modeling paradigm (Zheng et al. 2025; Yao et al. 2025) itself has also made significant progress. For instance, MAR (Li et al. 2024) uses masked multi-token prediction to improve quality and speed, while RandAR (Pang et al. 2025) enables flexible token ordering for new capabilities such as resolution extrapolation. To better capture spatial dependencies, models like VAR (Tian et al. 2024) and Infinity (Han et al. 2025) introduced coarse-to-fine prediction schemes. Compared to diffusion models, AR approaches generally offer superior prompt adherence and faster inference. However, their application to instruction-guided image editing remains underexplored. This challenging as it requires conditioning on both visual and text inputs, and our work aims to bridge this gap.
Figure 2: The overall architecture of VAREdit for instruction-guided image editing. VAREdit first encodes and quantizes the images into multi-scale residuals and maps the instructions into textual token embeddings. These features are organized as the finest-scale source feature $\mathbf{F}_{K}^{(src)}$, the pooled textual representation $\widetilde{\mathbf{F}}_{0}^{(tgt)}$, the coarse-to-fine target features $\widetilde{\mathbf{F}}_{1:K-1}^{(tgt)}$, and then sent to the VAR Transformer. The source feature $\mathbf{F}_{K}^{(src)}$ is further sent to the SAR module in the first self-attention layer to address the scale mismatch issue, while the textual token embeddings are also used for cross-attention calculations of key and value matrices. The ground truth residuals $\mathbf{R}_{1:K}$ guide the training of the last $K$ output residuals $\hat{\mathbf{R}}_{1:K}$. During inference, the residuals $\hat{\mathbf{R}}_{1:K}$ are predicted autoregressively, which are then cumulated and decoded to the edited image.
Methodology
@@
We first review the visual autoregressive (VAR) modeling paradigm. Then, we introduce VAREdit, a novel framework that reframes instruction-guided image editing as a multi-scale conditional generation task. Finally, we analyze the challenges of source image conditioning and present our Scale-Aligned Reference (SAR) module, a targeted solution to the scale-mismatch problem in naive conditioning.
Preliminary
@@
Visual autoregressive (VAR) model generally comprises a multi-scale visual tokenizer and a Transformer-based generation model. The process begins with an encoder $\mathcal{E}$ that maps an image $\mathbf{I}$ to a continuous feature representation $\mathbf{F}=\mathcal{E}(\mathbf{I})\in\mathbb{R}^{h\times w\times d}$. A quantizer $\mathcal{Q}$ then decomposes $\mathbf{F}$ into a hierarchy of $K$ discrete residual maps $\mathbf{R}_{1:K}=\mathcal{Q}(\mathbf{F})$. These maps follow a coarse-to-fine structure, where the spatial resolution $(h_{k},w_{k})$ of each residual $\mathbf{R}_{k}$ increases with its scale index $k$. Then, the Transformer model predicts the residuals in an autoregressive manner:
@@
Concretely, to predict the residual map for the next scale $\mathbf{R}_{k+1}$, the model first computes an intermediate feature representation $\mathbf{F}_{k}$, by aggregating all previously generated residuals $\mathbf{R}_{1:k}$:
@@
VAREdit
@@
We introduce VAREdit, a framework that reframes instruction-guided image editing as a conditional, multi-scale prediction problem. As illustrated in Figure 2, VAREdit leverages a pre-trained VAR model as its foundation and autoregressively generates the target residual maps $\mathbf{R}^{(tgt)}$ conditioned on a source image $\mathbf{I}^{(src)}$ and a textual instruction $\mathbf{t}$:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{R}_{1:k-1}^{(tgt)},\mathbf{I}^{(src)},\mathbf{t}).$ (3)
@@
A key design challenge is how to effectively and efficiently incorporate the source image $\mathbf{I}^{(src)}$ to guide the multi-scale generation process.
Vanilla Full-scale Condition
@@
A straightforward approach is to condition the generation on the complete multi-scale features of the source image, $\mathbf{F}_{1:K}^{(src)}$. This is achieved by prepending the sequence of all source tokens to the target sequence, which allows the model to reference any scale source residual when generating the target residual. The conditional likelihood becomes:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})$ $=p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{F}_{1:K}^{(src)},\mathbf{t})$ (4)
$=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{R}_{1:k-1}^{(tgt)},\mathbf{F}_{1:K}^{(src)},\mathbf{t}).$
@@
While this method provides a comprehensive, scale-by-scale reference for the editing task, it is computationally expensive. Doubling the sequence length results in a quadratic increase in the cost of self-attention, rendering it impractical for high-resolution editing. Moreover, providing multiple source scale features may bring redundant or conflicting information for predicting single scale target features, potentially degrading editing quality.
Finest-scale Condition
@@
To address the prohibitive cost of full-scale conditioning, we propose a more efficient strategy to condition solely on the finest-scale source feature, $\mathbf{F}_{K}^{(src)}$. This approach is motivated by the hierarchical nature of the visual tokenizer: the finest scale encapsulates the most detailed, high-frequency information from the source image, which is often the most critical for guiding an edit. This simplification slims down the likelihood to:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})$ $=p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{F}_{K}^{(src)},\mathbf{t})$ (5)
$=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{R}_{1:k-1}^{(tgt)},\mathbf{F}_{K}^{(src)},\mathbf{t}).$
@@
In this way, only the tokens from $\mathbf{F}_{K}^{(src)}$ are prepended to the target sequence. While this strategy dramatically reduces sequence length and thus alleviates the computational bottleneck of the vanilla Full setting, it introduces a critical scale mismatch challenge. The model is tasked with predicting coarse-grained target image structure while only having access to the fine-grained, local details in the source condition, which is insufficient for precise editing.
Scale Dependency Analysis

### v2

@@ 1 Introduction
Instruction-guided image editing (Brooks et al., 2023; Zhang et al., 2023; Sheynin et al., 2024) represents a significant advance in generative AI, shifting the paradigm from pure synthesis towards fine-grained interactive control. This progress has been propelled by two primary factors: the curation of large-scale, high-quality editing datasets (Ge et al., 2024; Ye et al., 2025) and the rise of diffusion models (Ho and others, 2020; Qian et al., 2024; Peebles and others, 2023; Labs, 2024; Qian et al., 2025; Wan et al., 2025) as the mainstream architecture. In particular, the impressive visual fidelity is largely derived from the iterative denoising process in diffusion models. However, this core mechanism also imposes a critical limitation. The global nature of the denoising process often leads to unintended edits, causing modifications to “bleed” into regions that should remain unchanged. Simultaneously, the multi-step denoising process incurs substantial computational cost and limits real-time applications.
@@
Figure 1: VAREdit achieves high-precision performance in instruction-guided image editing. It excels across diverse editing scenarios, including object-level modifications (addition, replacement, removal), attribute changes (material, text, posture, style, color) and complex compositional edits.
In contrast to diffusion models, autoregressive (AR) models (Esser and others, 2021; Sun et al., 2024; Pang et al., 2025; Yao et al., 2025; Zheng et al., 2025) offer a fundamentally different paradigm. Instead of holistic iterative refinement, they synthesize an image causally in a sequential token-by-token manner. This compositional generation process is a natural fit for image editing. Specifically, it provides a flexible mechanism for preserving unchanged regions while precisely modifying edited regions, addressing the entanglement problem of diffusion models. More encouragingly, the pioneering visual autoregressive (VAR) modeling (Tian et al., 2024) breaks the spatial structural degradation through the scale-by-scale prediction strategy, leading to high-quality image generation. While the VAR paradigm has shown competitive performance in image synthesis (Kumbong et al., 2025; Han et al., 2025), its potential for instruction-guided image editing remains largely unexplored. Early training-free attempts (Wang et al., 2025) repurpose text-to-image models but lack the task-specific knowledge required for robust editing, causing them to lag significantly behind diffusion-based methods (Hou and others, 2024; Zhao et al., 2024; Zhang et al., 2025b).
@@
We address this critical gap by introducing VAREdit, a novel tuning-based VAR framework for instruction-guided image editing. VAREdit reframes image editing as a next-scale prediction problem, which generates multi-scale target features to achieve precise image modification. A central challenge in this paradigm is how to effectively condition the model on the source image. A straightforward approach is to use full-scale source conditions, yet it is computationally inefficient due to the lengthy input sequences. While the finest-scale-only conditional setting provides sufficient information and eliminates the redundant input tokens, it creates a severe scale mismatch since high-frequency details may disrupt the prediction of coarse target features. To resolve this issue, we perform a systematic analysis of inter-scale dependencies and uncover a key insight: the model’s sensitivity to scale-aware conditioning is concentrated in its first self-attention layer. Accordingly, we propose the Scale-Aligned Reference (SAR) module, a mechanism that injects the scale-matched source information only in the first self-attention layer, while other layers still use the finest-scale condition. As demonstrated in Figure 1, this design enables VAREdit to achieve exceptionally precise editing across diverse challenging editing scenarios. Our contributions are three-fold:
@@
  • We introduce VAREdit, a novel tuning-based visual autoregressive model for instruction-guided image editing. To the best of our knowledge, we are among the first to introduce next-scale visual prediction into instruction-guided image editing.
@@
  • We identify a scale-mismatch issue in the finest-scale conditioned VAREdit and propose the SAR module as an effective solution.
@@
  • VAREdit attains a new record on standard editing benchmarks, outperforming leading diffusion models in both editing adherence and generation efficiency.
2 Related Work
2.1 Instruction-Guided Image Editing
@@
Instruction-guided image editing (Brooks et al., 2023; Zhao et al., 2024; Yu et al., 2025; Cai et al., 2025) seeks to modify an image according to user’s textual instruction, which differs from caption-based approaches (Hertz et al., 2022; Feng et al., 2025) that often resynthesize an entire image from a global description. This field is predominantly driven by diffusion models and seminal work InstructPix2Pix (Brooks et al., 2023) established a powerful paradigm by channel-wise concatenating source and target images. Subsequent research advanced this by curating larger-scale datasets with “expert” models (Zhang et al., 2023; Sheynin et al., 2024; Zhao et al., 2024; Ge et al., 2024; Ye et al., 2025) or exploring alternative conditioning, like spatial concatenation for large-motion edits (Zhang et al., 2025b; Cai et al., 2025). However, diffusion-based models often produce unintended modifications and suffer from slow iterative sampling. To address these limitations, recent studies have widely explored autoregressive (AR) models for instruction-guided image editing. Some approaches leverage the capabilities of Multimodal Large Language Models (MLLMs) in high-quality semantical alignment (Fu et al., 2024; Zhou et al., 2025) for further diffusion-based visual generation processes. Other methods employ AR-based frameworks directly for the image editing task, integrating knowledge distillation (Mu and others, 2025), mask-aware region conditioning (Wu et al., 2025b) or in-context reasoning for interactive manipulation (Lai et al., 2025). While these methods have achieved promising results, they still follow the vanilla next-token prediction paradigm with potential risks of structural degradation. Our work builds on the pioneering visual autoregressive (VAR) (Tian et al., 2024) architecture that follows the next-scale prediction paradigm, a tuning-based approach that surpasses recent approaches in both edit quality and efficiency.
2.2 Autoregressive Image Generation
@@
Autoregressive (AR) models, a mainstay in language tasks, have been extended to image synthesis. They operate by tokenizing an image into a discrete sequence and then training a Transformer model to predict the next token. Early works use vector quantization (VQ) for tokenization (Van Den Oord and others, 2017; Esser and others, 2021; Chang et al., 2022; Yu et al., 2022), while recent advances such as lookup-free quantization (Yu et al., 2023; Zhao and others, 2024) have improved performance with large vocabularies. The AR modeling paradigm itself has also made significant progress. For instance, MAR (Li et al., 2024) uses masked multi-token prediction to improve quality and speed, while RandAR (Pang et al., 2025) enables flexible token ordering for new capabilities such as resolution extrapolation. To better capture spatial dependencies, models like VAR (Tian et al., 2024) and Infinity (Han et al., 2025) introduced coarse-to-fine prediction schemes. As AR approaches are well-suited for language modeling, they demonstrate a significant advantage in instruction-following performance (Lai et al., 2025; Zhang et al., 2025a) for image generation. Meanwhile, recent advances like hybrid parallelization and KV-scale mitigate the efficiency bottleneck (Kumbong et al., 2025; Li et al., 2025). These innovations are enhancing both generation quality and efficiency, enabling AR models to emerge as strong competitors to diffusion models, particularly in tasks requiring high fidelity to complex, compositional prompts.
3 Methodology
@@
We first review the visual autoregressive (VAR) modeling paradigm. Then, we introduce VAREdit, a novel framework that reframes instruction-guided image editing as a multi-scale conditional generation task. Finally, we analyze the challenges of various source image conditioning strategies and present the inspiration and the design of our Scale-Aligned Reference (SAR) module, a targeted solution to the scale-mismatch problem in naive conditioning.
3.1 Preliminary
@@
Visual autoregressive (VAR) model generally comprises a multi-scale visual tokenizer and a Transformer-based generation model. The process begins with an encoder $\mathcal{E}$ that maps an image $\mathbf{I}$ to a continuous feature representation $\mathbf{F}=\mathcal{E}(\mathbf{I})\in\mathbb{R}^{h\times w\times d}$. A quantizer $\mathcal{Q}$ then decomposes $\mathbf{F}$ into a hierarchy of $K$ discrete residual maps $\mathbf{R}_{1:K}=\mathcal{Q}(\mathbf{F})$. These maps follow a coarse-to-fine structure, where the spatial resolution $(h_{k},w_{k})$ of each residual map $\mathbf{R}_{k}$ increases with its scale index $k$. Then, the Transformer model predicts the residuals in an autoregressive manner (where the scale features are flattened to obtain the sequential representation):
@@
Concretely, to predict the residual map for the next scale $\mathbf{R}_{k+1}$, the model first computes an intermediate feature representation $\mathbf{F}_{k}$ by aggregating all previously generated residuals $\mathbf{R}_{1:k}$:
@@
Figure 2: The overall architecture of VAREdit for instruction-guided image editing. VAREdit first encodes the images into multi-scale residuals by a shared vector quantizer and maps the instructions into textual token embeddings. These features are organized as the finest-scale source feature $\mathbf{F}_{K}^{(src)}$, the pooled textual representation $\widetilde{\mathbf{F}}_{0}^{(tgt)}$, the coarse-to-fine target features $\widetilde{\mathbf{F}}_{1:K-1}^{(tgt)}$, and then sent to the VAR Transformer. The source feature $\mathbf{F}_{K}^{(src)}$ is further sent to the Scale-Aligned Reference (SAR) module in the first self-attention layer, while the textual token embeddings are also used for cross-attention calculations of key and value matrices. The ground truth residuals $\mathbf{R}_{1:K}^{(tgt)}$ guide the training of the last $K$ output residuals $\hat{\mathbf{R}}_{1:K}^{(tgt)}$. During inference, the residuals $\hat{\mathbf{R}}_{1:K}^{(tgt)}$ are predicted autoregressively, which are then aggregated to $\hat{\mathbf{F}}_{K}^{(tgt)}$ and then decoded to the edited image.
3.2 VAREdit
@@
We introduce VAREdit, a framework that reframes instruction-guided image editing as a conditional, multi-scale prediction problem. Figure 2 presents the overall architecture of VAREdit, which leverages a pre-trained VAR model as its foundation and autoregressively generates the target residual maps $\mathbf{R}^{(tgt)}$ conditioned on a source image $\mathbf{I}^{(src)}$ and a textual instruction $\mathbf{t}$:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{R}_{1:k-1}^{(tgt)},\mathbf{I}^{(src)},\mathbf{t})=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{F}_{1:k-1}^{(tgt)},\mathbf{I}^{(src)},\mathbf{t}),$ (3)
where the second equation comes from the fact that $\mathbf{F}_{k}^{(tgt)}$ integrates all previous $k$ scales’ residual information $\mathbf{R}_{1:k}^{(tgt)}$ according to Equation 2 and we take the spatial-dimension aligned form $\widetilde{\mathbf{F}}_{k}^{(tgt)}$ of the aggregated features as input. All scale features are flattened to obtain a sequential representation.
3.2.1 Conditioning Source Images
@@
As for the specific design of VAREdit, the key challenge is how to effectively and efficiently incorporate the source image $\mathbf{I}^{(src)}$ to guide the multi-scale generation process.
@@
Full-scale Conditioning. A straightforward approach is to condition the generation on the complete multi-scale features of the source image, $\mathbf{F}_{1:K}^{(src)}$. This is achieved by prepending the sequence of all source tokens to the target sequence, which allows the model to reference any scale source residual when generating the target residual. The conditional likelihood becomes:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})=p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{F}_{1:K}^{(src)},\mathbf{t})=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{F}_{1:k-1}^{(tgt)},\mathbf{F}_{1:K}^{(src)},\mathbf{t}).$ (4)
@@
While this method provides a comprehensive, scale-by-scale reference for the editing task, it is computationally expensive. Doubling the sequence length results in a quadratic increase $O(n^{2})$ in the cost of self-attention, rendering it impractical for high-resolution editing. Moreover, providing multiple source scale features may introduce redundant or conflicting information for target feature prediction, potentially degrading editing quality.
@@
Finest-scale Conditioning. To address the prohibitive cost of full-scale conditioning, we propose a more efficient strategy to condition solely on the finest-scale source feature, $\mathbf{F}_{K}^{(src)}$. This approach is motivated by the hierarchical nature of the visual tokenizer: the finest scale encapsulates the most detailed, high-frequency information from the source image, which is often the most critical for guiding an edit. This simplification slims down the likelihood to:
$p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{I}^{(src)},\mathbf{t})=p(\mathbf{R}_{1:K}^{(tgt)}|\mathbf{F}_{K}^{(src)},\mathbf{t})=\prod_{k=1}^{K}p(\mathbf{R}_{k}^{(tgt)}|\mathbf{F}_{1:k-1}^{(tgt)},\mathbf{F}_{K}^{(src)},\mathbf{t}).$ (5)
@@
In this way, only the tokens from $\mathbf{F}_{K}^{(src)}$ are prepended to the target sequence. While this strategy dramatically reduces sequence length and thus alleviates the computational bottleneck of the vanilla full-scale conditional setting, it introduces a critical scale mismatch challenge. The model is tasked with predicting coarse-grained target image structure while having access to only the fine-grained, local details in the source condition, which is insufficient for precise editing.
3.2.2 Scale Dependency Analysis